OpenStudy (anonymous):

Write a formula for this sequence: 1/5, 3/5, 1, 1 2/5...

jhonyy9 (jhonyy9):

1/5 , 3/5 , 5/5 , 7/5 , ... add 2 to nominator

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

A.-1/5 + 2/5 n B.1/5 + 2/5 n C.1 2/5 - 2/5 n are my options

OpenStudy (anonymous):

Thanks for trying though

OpenStudy (anonymous):

(2n-1)/5

jhonyy9 (jhonyy9):

yes cihan you are right

jhonyy9 (jhonyy9):

for n natural

OpenStudy (anonymous):

of your three options, wouldn't A be the closest for the start? -1/5 + 2/5 = 1/5 1/5 + 2/5 = 3/5 3/5 + 2/5 =1 1 + 2/5 = 1 2/5 and so on.

OpenStudy (anonymous):

and so i would answer 'A'